For a d-electron, the orbital angular momentum is
a. √6 (h/2π)
b. √2 (h/2π)
c. (h/2π)
d. 2 (h/2π)

Grade:upto college level

2 Answers

Deepak Patra
9 years ago
The orbital angular momentum (L) = √(l (l+1)) h/2π = √6 h/2π (l = 2 for d-orbital)
Kushagra Madhukar
3 years ago
Dear student,
Please find the attached solution to your solution.
 
The orbital angular momentum (L) = √(l (l+1)) h/2π = √6h/2π             (l = 2 for d-orbital)
Hence option a) is correct.
 
Hope it helps.
Thanks and regards,
Kushagra